OpenStudy (anonymous):

Simplify the expression (7x2 − 3x) + 4x2

OpenStudy (anonymous):

A. 8x B. 8x2 C. 3x2 − 3x D. 11x2 − 3x

OpenStudy (anonymous):

(7x2−3x)+4x2 =7z2−3x+4x2 =7x2+4x2−3x =11x2−3x

OpenStudy (anonymous):

I mean =7x2−3x+4x2 not z lol

OpenStudy (anonymous):

haha thanks